# Spectral Concentration at the Edge of Stability: Information Geometry of Kernel Associative Memory

### v1

@@
Kernel Hopfield Network, Information Geometry, Edge of Stability, Spectral Concentration, Self-Organized Criticality
@@
High-capacity kernel Hopfield networks exhibit a “Ridge of Optimization” characterized by extreme stability. While previously linked to ”Spectral Concentration,” its origin remains elusive. Here, we analyze the network dynamics on a statistical manifold, revealing that the Ridge corresponds to the “Edge of Stability,” a critical boundary where the Fisher Information Matrix becomes singular. We demonstrate that the apparent Euclidean force antagonism is a manifestation of Dual Equilibrium in the Riemannian space. This unifies learning dynamics and capacity via the Minimum Description Length principle, offering a geometric theory of self-organized criticality.
@@ 1 Introduction
Associative memory is fundamentally a geometric problem: how to embed discrete patterns as stable fixed points within a continuous state space. Recent advances using Kernel Logistic Regression (KLR) have demonstrated that learning can sculpt these landscapes to achieve capacities far exceeding classical limits [1, 2, 3]. Our previous phenomenological analysis identified a “Ridge of Optimization” where stability is maximized via a mechanism we termed Spectral Concentration—a state where the weight spectrum exhibits a sharp hierarchy [4].
@@
    The “Ridge” is the locus where the Fisher Information Matrix (FIM) undergoes a spectral phase transition, effectively minimizing the information dimension of the memory.
@@ 2.1 Statistical Manifold of Kernel Memory
Let $\mathcal{H}$ be a reproducing kernel Hilbert space with kernel $K(\cdot,\cdot)$. The state of the network is determined by the dual variables $\bm{\alpha}\in\mathbb{R}^{P\times N}$, where $P$ is the number of stored patterns and $N$ is the number of neurons. For a single neuron $i$, the probability of being in state $s_{i}=+1$ given the input pattern $\bm{\xi}^{\mu}$ is modeled by the logistic sigmoid function:
@@ 2.2 Fisher Information Matrix (FIM)
Our central hypothesis is that the “Ridge” phenomena emerge from the interplay between these two vector fields—the Euclidean force $\nabla L$ and the geometric curvature $G$.
@@ 3 Spectral Concentration as Information Compression
Our previous phenomenological study identified “Spectral Concentration” in the weight matrix $\bm{\alpha}$ as the key to stability. Here, we show that this is a direct consequence of the information geometry of the learned manifold.
@@ 3.1 Effective Dimensionality of Memory
Observation 1 (L-Shaped Spectrum): Numerical analysis reveals Fig. 1 reveals a striking contrast in the spectral structures. In the local regime (blue dashed line), the spectrum is flat ($\lambda_{k}\approx\lambda_{1}$), indicating a diffuse distribution of information where no dominant direction exists. In contrast, on the Ridge (red solid line), the spectrum exhibits a characteristic “L-shape”: the normalized eigenvalues drop by orders of magnitude immediately after the leading mode ($\lambda_{2}\ll\lambda_{1}$), yet the tail remains non-zero and flat ($\lambda_{k>1}>0$). This confirms that the Ridge achieves stability via Spectral Concentration—compressing noise into a single dominant mode while preserving the dimensionality required for memory capacity. Quantitatively, while the local regime exhibits $d_{\text{eff}}(G)\approx P$ (full utilization of dimensions), the Ridge collapses this to $d_{\text{eff}}(G)\approx 1.5\sim 5.0$, depending on the load. This drastic reduction confirms that the network performs optimal compression, retaining only the essential degrees of freedom.
@@ 4.1 Maximal Curvature at the Ridge
Figure 2: Phase diagram of the maximum eigenvalue of the Fisher Information Matrix, $\log_{10}(\lambda_{\max}(G))$. The Ridge (bright diagonal band) corresponds to the region where the manifold curvature is locally maximized ($\lambda_{\max}\gg 1$), effectively creating the “Edge of Stability.” In contrast, the dark region (bottom right) indicates information collapse due to saturation of the sigmoid function ($p\to 0/1$), leading to vanishing curvature.
Observation 2 (Criticality): Our phase diagram analysis (Fig. 2) reveals that $\lambda_{\max}(G)$ is maximized exactly on the Ridge. While in the local regime (top region), the curvature is relatively small ($\log_{10}\lambda_{\max}\approx-1.2$), on the Ridge, it reaches values orders of magnitude higher ($\log_{10}\lambda_{\max}\approx 1.5$), indicating the formation of extremely sharp attractors. Crucially, immediately beyond the Ridge (in the global regime, bottom right), the value drops precipitously ($\log_{10}\lambda_{\max}<-8.0$). This corresponds to a phase where the learning saturates ($p_{\mu}\to 0$ or $1$), causing the Fisher information to vanish due to the vanishing gradients of the sigmoid function. Thus, the Ridge represents the critical boundary where the system maximizes geometric curvature to the limit allowed before information collapse occurs.
@@ 4.2 The Cliff of Optimization
Thus, the “Ridge of Optimization” can be geometrically defined as the Edge of Stability: the manifold boundary where the curvature is maximal.
@@ 5 Dual Geometry and Force Antagonism
Our previous work identified a phenomenological “Force Antagonism” on the Ridge: the Direct Force $\bm{F}_{d}$ and Indirect Force $\bm{F}_{i}$ become massive and strongly anti-correlated ($\bm{F}_{d}\approx-\bm{F}_{i}$). We now reveal the geometric origin of this antagonism.
@@
Figure 3: Comparison of Gradient Norms: (a) Euclidean Gradient $\|\nabla L\|^{2}$ vs. (b) Riemannian (Natural) Gradient $\|\tilde{\nabla}L\|_{G^{-1}}^{2}$. On the Ridge, the Euclidean norm diverges (bright band in a), reflecting the massive “Force Antagonism.” However, the Riemannian norm is minimized (dark band in b), revealing that the system is actually in a state of Dual Equilibrium with respect to the intrinsic geometry. This confirms that the apparent instability in parameter space is resolved by the curvature of the statistical manifold.
@@ 5.2 Geometric Derivation of Antagonism
We define the “Indirect Force” in the language of information geometry as the correction term arising from the curvature of the manifold. The Natural Gradient relation $\nabla L=G\tilde{\nabla}L$ implies that the observed Euclidean force is the intrinsic force distorted by the metric $G$. On the Ridge, we found that the spectrum of $G$ is dominated by a leading eigenvalue $\lambda_{1}\gg 1$ with eigenvector $\bm{v}_{1}$. Approximating $G\approx\lambda_{1}\bm{v}_{1}\bm{v}_{1}^{\top}$, the relationship becomes:
@@ 6 Discussion
In this study, we have reinterpreted the dynamics of KLR-trained Hopfield networks through the lens of Information Geometry. Our analysis reveals that the phenomenological “Ridge of Optimization” is not merely a sweet spot of hyperparameters, but a manifestation of fundamental geometric principles: Spectral Concentration, Edge of Stability, and Dual Equilibrium.
@@ 6.1 Self-Organized Criticality in Memory
Our finding that the optimal memory state lies at the brink of information collapse (Fig. 2) aligns with the hypothesis of Self-Organized Criticality (SOC) [9] in biological neural networks. The brain is thought to operate at a critical point to maximize information transmission and dynamic range. Similarly, our KLR Hopfield network tunes itself to the “Edge of Stability,” maximizing the curvature of the statistical manifold to create deep attractors, while stopping just short of the singular region where memory capacity vanishes. This suggests that SOC may be a universal requirement for high-capacity associative memory.
@@ 6.3 The Geometry of Attention Mechanisms
Modern Hopfield Networks are mathematically equivalent to the attention mechanism in Transformers [11]. Our geometric framework implies that the effectiveness of attention layers may also rely on specific spectral structures in their Key-Query matrices. If attention heads also operate at an “Edge of Stability” to sharpen their focus, our duality theory (Euclidean force vs. Riemannian curvature) could provide a new tool for analyzing and stabilizing the training of large-scale Transformer models. Future work will investigate whether trained Transformers exhibit similar spectral signatures to those found on the Ridge.
@@ 7 Conclusion
This geometric framework unifies the apparently contradictory observations of instability (high curvature, Edge of Stability) and stability (deep attractors, Dual Balance). The “Ridge of Optimization” is not merely a parameter sweet spot, but a manifestation of Self-Organized Criticality in the information geometry of neural networks. The network tunes itself to the brink of instability, maximizing curvature to carve the deepest possible attractor basins, while maintaining dual equilibrium to ensure learning convergence.

### v2

@@
Hopfield Network, Information Geometry, Edge of Stability, Spectral Concentration, Self-Organized Criticality
@@
High-capacity kernel Hopfield networks exhibit a Ridge of Optimization characterized by extreme stability. While previously linked to Spectral Concentration, its origin remains elusive. Here, we analyze the network dynamics on a statistical manifold, revealing that the Ridge corresponds to the Edge of Stability, a critical boundary where the Fisher Information Matrix becomes singular. We demonstrate that the apparent Euclidean force antagonism is a manifestation of Dual Equilibrium in the Riemannian space. This unifies learning dynamics and capacity via the Minimum Description Length principle, offering a geometric theory of self-organized criticality.
@@ 1 Introduction
Associative memory is fundamentally a geometric problem: how to embed discrete patterns as stable fixed points within a continuous state space. Recent advances using Kernel Logistic Regression (KLR) have demonstrated that learning can sculpt these landscapes to achieve capacities far exceeding classical limits [1, 2, 3]. Our previous phenomenological analysis identified a Ridge of Optimization where stability is maximized via a mechanism we termed Spectral Concentration, defined as a state where the weight spectrum exhibits a sharp hierarchy [4].
@@
    The Ridge is the locus where the Fisher Information Matrix (FIM) undergoes a spectral phase transition, effectively minimizing the information dimension of the memory.
@@ 2.1 Statistical Manifold of Kernel Memory
Let $\mathcal{H}$ be a reproducing kernel Hilbert space with kernel $K(\cdot,\cdot)$. The state of the network is determined by the dual variables $\bm{\alpha}\in\mathbb{R}^{P\times N}$, where $P$ is the number of stored patterns and $N$ is the number of neurons. For a single neuron $i$, the probability of being in state $s_{i}=+1$ given the input pattern $\bm{\xi}^{\mu}$ is modeled by the logistic sigmoid function $\sigma(\cdot)$:
@@ 2.2 Fisher Information Matrix (FIM)
Our central hypothesis is that the Ridge phenomena emerge from the interplay between these two vector fields—the Euclidean force $\nabla L$ and the geometric curvature $G$.
@@ 3 Spectral Concentration as Information Compression
Our previous phenomenological study identified Spectral Concentration in the weight matrix $\bm{\alpha}$ as the key to stability [4]. Here, we show that this is a direct consequence of the information geometry of the learned manifold.
@@ 3.1 Effective Dimensionality of Memory
Observation 1 (L-Shaped Spectrum): Numerical analysis reveals Fig. 1 reveals a striking contrast in the spectral structures. In the local regime (blue dashed line), the spectrum is flat ($\lambda_{k}\approx\lambda_{1}$), indicating a diffuse distribution of information where no dominant direction exists. In contrast, on the Ridge (red solid line), the spectrum exhibits a characteristic “L-shape”: the normalized eigenvalues drop by orders of magnitude immediately after the leading mode ($\lambda_{2}\ll\lambda_{1}$), yet the tail remains non-zero and flat ($\lambda_{k>1}>0$). This confirms that the Ridge achieves stability via Spectral Concentration, effectively compressing noise into a single dominant mode while preserving the dimensionality required for memory capacity. Quantitatively, while the local regime exhibits $d_{\text{eff}}(G)\approx P$ (full utilization of dimensions), the Ridge collapses this to $d_{\text{eff}}(G)\approx 1.5\sim 5.0$, depending on the load. This drastic reduction confirms that the network performs optimal compression, retaining only the essential degrees of freedom.
@@ 4.1 Maximal Curvature at the Ridge
Figure 2: Phase diagram of the maximum eigenvalue of the Fisher Information Matrix, $\log_{10}(\lambda_{\max}(G))$ (averaged over 5 trials). The Ridge (bright diagonal band) corresponds to the region where the manifold curvature is locally maximized ($\lambda_{\max}\gg 1$), effectively creating the “Edge of Stability.” In contrast, the dark region (bottom right) indicates information collapse due to saturation of the sigmoid function ($p\to 0/1$), leading to vanishing curvature.
Observation 2 (Criticality): Our phase diagram analysis (Fig. 2) reveals that $\lambda_{\max}(G)$ is maximized exactly on the Ridge. While in the local regime (top region), the curvature is relatively small ($\log_{10}\lambda_{\max}\approx-1.2$), on the Ridge, it reaches significantly higher values ($\log_{10}\lambda_{\max}>1.0$), indicating the formation of extremely sharp attractors. Crucially, immediately beyond the Ridge (in the global regime, bottom right), the value drops precipitously ($\log_{10}\lambda_{\max}<-6.0$). This corresponds to a phase where the learning saturates ($p_{\mu}\to 0$ or $1$), causing the Fisher information to vanish due to the vanishing gradients of the sigmoid function. Thus, the Ridge represents the critical boundary where the system maximizes geometric curvature to the limit allowed before information collapse occurs.
@@ 4.2 The Cliff of Optimization
Thus, the Ridge of Optimization can be geometrically defined as the Edge of Stability: the manifold boundary where the curvature is maximal.
@@ 5 Dual Geometry and Force Antagonism
Our previous work identified a phenomenological Force Antagonism on the Ridge: the Direct Force $\bm{F}_{d}$ and Indirect Force $\bm{F}_{i}$ become massive and strongly anti-correlated ($\bm{F}_{d}\approx-\bm{F}_{i}$). We now reveal the geometric origin of this antagonism.
@@
Figure 3: Comparison of Gradient Norms: (a) Euclidean Gradient $\|\nabla L\|^{2}$ vs. (b) Riemannian (Natural) Gradient $\|\tilde{\nabla}L\|_{G^{-1}}^{2}$ (averaged over 5 trials). On the Ridge, the Euclidean norm diverges (bright band in a), reflecting the massive “Force Antagonism.” However, the Riemannian norm is minimized (dark band in b), revealing that the system is actually in a state of Dual Equilibrium with respect to the intrinsic geometry. This confirms that the apparent instability in parameter space is resolved by the curvature of the statistical manifold.
@@ 5.2 Geometric Derivation of Antagonism
We define the Indirect Force in the language of information geometry as the correction term arising from the curvature of the manifold. The Natural Gradient relation $\nabla L=G\tilde{\nabla}L$ implies that the observed Euclidean force is the intrinsic force distorted by the metric $G$. On the Ridge, we found that the spectrum of $G$ is dominated by a leading eigenvalue $\lambda_{1}\gg 1$ with eigenvector $\bm{v}_{1}$. Approximating $G\approx\lambda_{1}\bm{v}_{1}\bm{v}_{1}^{\top}$, the relationship becomes:
@@ 6 Discussion
In this study, we have reinterpreted the dynamics of KLR-trained Hopfield networks through the lens of Information Geometry. Our analysis reveals that the phenomenological Ridge of Optimization is not merely a sweet spot of hyperparameters, but a manifestation of fundamental geometric principles: Spectral Concentration, Edge of Stability, and Dual Equilibrium.
@@ 6.1 Self-Organized Criticality in Memory
Our finding that the optimal memory state lies at the brink of information collapse (Fig. 2) aligns with the hypothesis of Self-Organized Criticality (SOC) [9] in biological neural networks. The brain is thought to operate at a critical point to maximize information transmission and dynamic range. Similarly, our KLR Hopfield network tunes itself to the Edge of Stability, maximizing the curvature of the statistical manifold to create deep attractors, while stopping just short of the singular region where memory capacity vanishes. This suggests that SOC may be a universal requirement for high-capacity associative memory.
@@ 6.3 The Geometry of Attention Mechanisms
Modern Hopfield Networks are mathematically equivalent to the attention mechanism in Transformers [11]. Our geometric framework implies that the effectiveness of attention layers may also rely on specific spectral structures in their Key-Query matrices. If attention heads also operate at an Edge of Stability to sharpen their focus, our duality theory (Euclidean force vs. Riemannian curvature) could provide a new tool for analyzing and stabilizing the training of large-scale Transformer models. Future work will investigate whether trained Transformers exhibit similar spectral signatures to those found on the Ridge.
@@ 7 Conclusion
This geometric framework unifies the apparently contradictory observations of instability (high curvature, Edge of Stability) and stability (deep attractors, Dual Balance). The Ridge of Optimization is not merely a parameter sweet spot, but a manifestation of Self-Organized Criticality in the information geometry of neural networks. The network tunes itself to the brink of instability, maximizing curvature to carve the deepest possible attractor basins, while maintaining dual equilibrium to ensure learning convergence.